3. A metal banana-shaped earring has a net charge of -2.0 µC while a metal apple-shaped earring has a net charge of 8.0 uC. The earrings are 3.0 m apart. (A) What is the electric force between the two earrings? Is it attractive or repulsive? (B) The two earrings are brought together, touching one another for a few seconds, after which the earrings are once again placed 3.0 m apart. What is the electric force between the two earrings now? Is it attractive or repulsive?


1
Expert's answer
2021-10-31T18:15:48-0400

1)

"F_1=\frac{k|q_1||q_2|}{r^2}=16~mN,"

"q_1q_2<0" – attractive,

2)

"q=\frac{q_1+q_2}2=3~\mu C,"

"F_2=\frac{kq^2}{r^2}=9~mN,"

"q^2>0" –repulsive.